Dell Camping Equipment, Inc. just issued a 10 percent, pays interest semiannually that matures in 25 years. Investors require 12 percent return for bearing risk of investing in this bond. (5)
i. How much maximum the investors will be willing to pay to purchase Dell bonds?
ii. Calculate the yield-to-maturity of Dell Bondholders assuming the bonds are selling at equilibrium price
1
Expert's answer
2020-05-27T16:18:40-0400

i) P-market bond price

D-dividens

N-nominal price

"P=D*\frac{(1/1.06)^{50}-1}{1/1.06-1}+\frac{N}{\frac{1}{1.06}^{50}}"

"P=D*16.7+N*18.42"

"D=0.1*N"

"P=N*20.02"

ii) P=N

"N=0.1*N*16.7+N*18.42"

"1=20.02"

One bond will give in 20.02 times more than cost